OpenStudy (anonymous):

FIND THE DERIVATIVE OF y^2= b^2/2x+b at O,b please help me.

OpenStudy (anonymous):

Is it: \[\large y^2 = \frac{b^2}{2x} + b\]

OpenStudy (anonymous):

b is in the denominator with 2x

OpenStudy (anonymous):

\[\large y^2 = \frac{b^2}{2x+b} \]

OpenStudy (anonymous):

Take the derivative implicitly..

OpenStudy (helder_edwin):

do what @waterineyes said do you know this?

OpenStudy (anonymous):

yes it is implicit differentiation.

OpenStudy (anonymous):

yy'=-b^2/(2x+b)^2

OpenStudy (anonymous):

Implicitly means: \[2y.dy = (b^2(-1)(2x+b)^{-2} \times 2).dx\] \[\large y.dy = (-b^2(2x + b)^{-2}).dx\] \[\large \frac{dy}{dx} = \frac{-b^2}{y(2x + b)^2}\] Now put x = 0 and y = b here..

OpenStudy (anonymous):

thanks @waterineyes! ((=

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

@waterineyes wait where did you get -1?

OpenStudy (anonymous):

Tell me the derivative of \(x^{-2}\) ??

OpenStudy (anonymous):

-2x^-3

OpenStudy (anonymous):

Yes I have done the same there..

OpenStudy (anonymous):

I have written: \[\frac{1}{2x+b} \quad as \quad (2x+b)^{-1}\] Now take its derivative.. Getting??

OpenStudy (anonymous):

oooh i get it. thanks a lot

OpenStudy (anonymous):

\[\frac{d}{dx}(2x + b)^{-1} = (-1) \times (2x + b)^{-2} \times 2\]

OpenStudy (anonymous):

Thanks! =]]

OpenStudy (anonymous):

Welcome buddy..